If P(A)=0.7,P(B)=0.1, and A and B are mutually exclusive, find P(A or B).
Is the answer 0?

Answer by ikleyn(52781)

.

If  A and B are mutually exclusive, it means that the events A and B are DISJOINT,  i.e.  CAN  NOT  occur simultaneously. 


Hence,


    P(A or B) = P(A) + P(B) = 0.7 + 0.1 = 0.8.


Solved, explained and completed.


Learn the basics of the Probability theory (!)